# Learning-to-Defer with Expert-Conditioned Advice

### 15 Mar 2026 v1

@@ C.5 CLIP Prompt-Escalation Experimental Details
Policies are trained on frozen CLIP-B/32 image embeddings of dimension $512$, using the same augmented surrogate at $\tau=1$ as in the other experiments. For every image, every expert, and every prompt family, we precompute the expert prediction and the realized cost, so all methods are trained and evaluated on exactly the same expert-advice outcome tensor. The augmented policy, L2D, best-fixed, partial-randomization, and joint-random baselines are exactly those defined in Section C.1, instantiated on this CLIP expert pool and prompt-family advice set.
@@
The tables below mirror those of the previous subsections. Table  15 first shows how each CLIP expert’s preferred advice action changes with the advice cost. Table  16 then evaluates full policies on the true deployment metric, and the detailed tables at $\lambda=0$, $\lambda=0.15$, and $\lambda=10$ make the cost-dependent expert-advice optimum explicit.
@@
Table 16 shows a clear cost-dependent transition. Our method attains the lowest mean true loss throughout the low- and moderate-cost regime, improving on both L2D and the best fixed pair. As $\lambda$ increases, the gap narrows, which is expected because the optimal behavior approaches the no-advice regime. At $\lambda=10$, our method is slightly above L2D in mean loss, consistent with the fact that querying has effectively become too expensive to justify. The partial-randomization baselines support the same conclusion as in the other experiments: the gain comes from coordinating expert choice and advice choice rather than from either component in isolation.

### 19 Mar 2026 v2

@@ C.5 CLIP Prompt-Escalation Experimental Details
Input features are frozen CLIP-B/32 image embeddings (dimension $512$), standardized with z-score normalization. All learned methods — our augmented surrogate and the L2D baseline — share the same backbone: a two-layer MLP with hidden dimensions $(512,256)$, ReLU activations, and no dropout. They are trained with AdamW (weight decay $10^{-3}$, gradient clipping at norm $10$) for $400$ epochs with batch size $256$ and a cosine learning-rate scheduler with $5\%$ linear warmup. We use a learning rate of $10^{-2}$ for L2D, and $10^{-3}$ for our approach. The augmented comp-sum surrogate uses $\tau=1$; the L2D baseline uses the same architecture but is restricted to the k=0k{=}0 slice and trained with the standard cross-entropy deferral surrogate. For every image, every expert, and every prompt family, we precompute the expert prediction and the realized cost, so all methods are trained and evaluated on exactly the same expert-advice outcome tensor. The best-fixed, partial-randomization, and joint-random baselines are exactly those defined in Section C.1, instantiated on this CLIP expert pool and prompt-family advice set.
@@
The tables below mirror those of the previous subsections. Table 15 first shows how each CLIP expert’s preferred advice action changes with the advice cost. Table 16 then evaluates full policies on the true deployment metric, and the detailed tables at $\lambda=0$, $\lambda=0.15$, and $\lambda=10$ make the cost-dependent expert-advice optimum explicit.
@@
Table 16 shows a clear cost-dependent transition. Our method attains the lowest mean true loss throughout the low- and moderate-cost regime, improving on both L2D and the best fixed pair. As $\lambda$ increases, the gap narrows, which is expected because the optimal behavior approaches the no-advice regime. At $\lambda=10$, our method ties L2D, consistent with the fact that querying has become too expensive to justify and the learned policy ceases to acquire advice. The partial-randomization baselines support the same conclusion as in the other experiments: the gain comes from coordinating expert choice and advice choice rather than from either component in isolation.